OpenStudy (anonymous):

what will the population of a city be if it currently has a population of 15,000 and grows at a rate of 2.5% per year for 5 years?

OpenStudy (anonymous):

16875

OpenStudy (anonymous):

population = 15000(1+2.5%)^5 = 15000(1.025)^5 = 16971

OpenStudy (aravindg):

16971